# Meta Engine: A Unified Semantic Query Engine on Heterogeneous LLM-Based Query Systems

### v1

@@ Abstract.
With the increasingly use of multi-modal data, semantic query has become more and more demanded in data management systems, which is an important way to access and analyze multi-modal data. As unstructured data, most information of multi-modal data (text, image, video, etc) hides in the semantics, which cannot be accessed by the traditional database queries like SQL. Comparing to traditional queries, semantic queries involve natural language to process the multi-modal data based on their semantics. Given the power of Large Language Model (LLM) in understanding semantics and processing natural language, in recent years several LLM-based semantic query systems have been proposed, to support semantic querying over unstructured data. However, this rapid growth has produced a fragmented ecosystem. Applications face significant integration challenges due to (1) disparate APIs of different semantic query systems and (2) a fundamental trade-off between specialization and generality. Many semantic query systems are highly specialized (e.g., DocETL for long documents, StructGPT for table QA), offering state-of-the-art performance within a single modality but struggling with multi-modal data. Conversely, some ”all-in-one” systems (e.g., LOTUS) handle multiple modalities but often exhibit suboptimal performance compared to their specialized counterparts in specific modalities. In short, the existing systems still have significant limitations when applied to complex multi-modal tasks.
This paper introduces Meta Engine, a novel ”query system on query systems”, designed to resolve those aforementioned challenges. Meta Engine is a unified semantic query engine that integrates heterogeneous, specialized LLM-based query systems. It combines the specialities of different systems to handle complex scenarios, and provides a unified natural language interface. Its architecture comprises five key components: (1) a Natural Language (NL) Query Parser to accept and break down complex NL queries into simpler sub-queries, (2) an Operator Generator that generates candidate operators for each sub-query, (3) a Query Router to dispatch sub-queries to the most appropriate specialized systems, (4) a set of Adapters to abstract and unify the heterogeneous APIs of underlying systems, and (5) a Result Aggregator to merge, deduplicate, and synthesize a coherent final answer. This design enables Meta Engine to combine the distinct strengths of different semantic query systems, achieving high-performance multi-modal semantic processing from a unified entry point. In the evaluation, Meta Engine consistently outperforms all baselines, yielding 3–6x higher F1 in most cases and up to  24x on specific datasets.
@@ 1. Introduction
The advance of Large Language Models (LLMs) leads to a paradigm shift in data analytics, resulting in the development of advanced semantic query systems. (Cappuzzo et al., 2020; Toenshoff et al., 2023; Bandyopadhyay et al., 2020; Patel et al., 2024; Shankar et al., 2024; Jiang et al., 2023; Khattab et al., 2023; Zhang et al., 2025). They enable applications to semantically reason and analyze over complex (like multi-modal) datasets using natural language.
Despite this rapid progress, the current ecosystem of LLM-based query systems is highly fragmented, presenting critical challenges that hinder the comprehensive use of them on complex tasks. First, the different tools result in disparate and incompatible APIs. Usually each specific system requires a unique design due to their specific scenarios. Consequently, developers seeking to build applications that leverage the strengths of multiple systems are forced to write complex integration logic, increasing both development and maintenance overhead. Second challenge is a crucial trade-off between specialization and generality. Many state-of-the-art semantic query systems are highly specialized, achieving exceptional performance by optimizing for a single modality. For instance, DocETL (Shankar et al., 2024) excels at processing large-scale long text documents, while StructGPT (Jiang et al., 2023) is designed specifically for semantic querying of tabular data. Their high specialization in specific domain/modality prevents them from integrating and processing information across different modalities. Conversely, ”all-in-one” systems like LOTUS (Patel et al., 2024) attempt to provide a unified solution for multi-modal data (e.g., table, text, and image), but the cost of such a generalization is often suboptimal performance on specific single modality, like the image QA quality of LOTUS is usually lower than a pure VLM.
@@
In this paper, we propose Meta Engine, a novel ”query system over query systems”. Meta Engine functions as an comprehensive framework, integrating heterogeneous LLM-based semantic query systems into a cohesive pipeline. By adaptively scheduling and calling the different systems, it can intelligently combine their specialities to achieve a significant performance gain. Additionally, Meta Engine provides a unified natural language (NL) querying interface, keeping the disparate APIs of underlying systems transparent to users. Note that although with an NL interface that allows users to query with NL questions, Meta Engine is not a question-answering (QA) system, but a general semantic query system. Meta Engine supports not only typical question answering tasks, but also broader data analytic tasks. In summary, the task of Meta Engine is modeled as such: given a supporting document repository with documents in different modalities (text, image and table), Meta Engine will process users’ queries and return the results. For example, in one of our evaluation dataset  (Rahman et al., 2025), each natural language query specifies an analytical task over structured tables, text and images. An example query and supporting documents are shown in Figure 1.
@@
Meta Engine includes several major components: (1) Natural Language Query Parser: This component accepts the user queries in natural language and intelligently preprocesses them. This mainly includes two modules, query complexity checker and query decomposer. The complexity checker determines whether the query is complex enough, and for a complex query (e.g., multi-hop queries), the decomposer will split it into a set of simpler sub-queries, each corresponding to a specific task and data modality, like the example shown in Figure 5.
Furthermore, we propose a novel hybrid sub-query generation method: initial decomposition + iterative sub-query refinement. Specifically, the query is first decomposed into sub-queries in one step, then the sub-queries are executed one by one while each sub-query is refined with the execution results from the proceeding sub-query if they have dependency. By such a dynamic refinement, the sub-queries are stated more clearly and the system avoids wasting time in duplicate computation across sub-queries.
(2) Operator Generator:
After decomposition, Meta Engine analyzes each sub-query’s intent and required evidence type, and maps it to a set of candidate operators using a confidence-based ranking mechanism rather than selecting a single operator. Specifically, the generator predicts a ranked list of operators, sorted by confidence scores, where higher-ranked operators are considered more likely to successfully answer the sub-query. To make routing efficient, this step relies only on lightweight metadata rather than full document contents.
(3) Query Router: The operators are then passed to the query router. The query router intelligently routes the opeartor to the most appropriate underlying specialist system (e.g., routing the text processing operator to DSPy and the image information extraction operator to VLM). Note that each operator could have multiple choices, like LOTUS, DSPy and VLM can all execute image operator. We design two types of routers: A learned router which is a query-dependent classifier on top of a pre-trained LLM, and a statistic router that is query-agnostic, based on history statistics. Our evaluation shows the learned router is overall better, but statistic router outperforms it in some cases.
(4) Adapters: Adapters serve as wrappers for the underlying LLM query systems. Each adapter conforms to a standardized internal interface while managing the unique API, request format, and authentication specifics of the system it wraps. This abstracts away the complexity and heterogeneity of the individual systems.
(5) Result Aggregator: As the adapters return results for their respective sub-queries, this final component aggregates the individual pieces of information and generates the final results. It refines the outputs (e.g., making long answers concise, extracting key answer keywords) and then merges and deduplicates (if necessary) the results from the different systems to synthesize a single, comprehensive, and accurate answer for the user.
By orchestrating the execution across specialized systems, Meta Engine resolves the specialist-generalist trade-off, enabling applications to benefit from high-performance, multi-modal processing through a single, unified natural language API.
@@
    We propose the Meta Engine architecture, a ”query system on query systems” that integrates diverse, specialized LLM query systems.
@@
    We introduce several components (Parser, Operator Generator, Router, Adapters, Aggregator) and novel optimization techniques to manage query decomposition, intelligent routing, execution and result synthesis, effectively combining the specialities of heterogeneous LLM query systems.
@@
    We integrate a unified natural language interface that abstracts the complexity of underlying APIs, simplifying the development of high-performance, multi-modal semantic applications on top of our system.
@@
    We conduct extensive evaluation that shows the effectiveness of our proposed system. We also perform ablation study to show the impacts of each major components in the system.
The remainder of this paper is structured as follows: Section 2 introduces prior works related to Meta Engine. Section 3 and 4 introduce the architecture, workflow and technical details of our method. Section 5 is the experiment section, analyzing the end-to-end results from different methods and the ablation studies for our method.
@@ 3. Architecture
Figure 2. Architecture and End-to-End Execution Pipeline of Meta Engine for Multimodal Queries
Figure 2 presents the architecture and end-to-end workflow of Meta Engine to process a query.
@@ 3.1. Architecture
Meta Engine is designed as a modular system that coordinates the processing of complex, multimodal queries through a series of specialized components. The process initiates at the Query Parser (shown in Figure 2 Stage 1), which acts as the initial interface for user queries. This component functions as a complexity checker and query decomposer, inputting a query and outputting one or more sub-queries.
The Operator Generator (shown in Figure 2 Stage 2) is responsible for identifying and ranking the most appropriate analytical operators for each sub-query. It determines the optimal operator for each sub-query with the help of Query Router and Adapters.
The execution layer is comprised of the Query Router and various Adapters, such as those for LOTUS, DocETL, and DSPy which wrap existing semantic query systems. These adapters interface directly with the multi-modal data sources (tables, text, and images) to perform the required reasoning. Router takes charge of sending each sub-query to the proper adapter to execute.
Finally, the partial results from these isolated sub-query executions are synthesized by the Final Aggregator (shown in Figure 2 Stage 3). This component aggregates the answers from all sub-queries, filtering out noise and refining the output to produce a concise, accurate final answer for the user.
Meta Engine demonstrates a significant improvement in result quality compared to the state-of-the-art semantic query systems. By decomposing complex queries and routing them to specialized underlying systems, Meta Engine consistently achieves higher accuracy in our evaluation.
A core advantage of the architecture is its modular design, specifically the use of Adapters to wrap underlying LLM systems. This design decouples the reasoning logic from the backend execution, allowing for the seamless addition or removal of underlying systems (e.g., LOTUS, DSPy, DocETL) without overhauling the core engine. With new powerful systems emerge, they can be easily integrated as new adapters.
3.2. Workflow
The system will first check and potentially decompose the original user query, as shown in Figure 2 stage 1:
  1. (1)
    Query complexity checker determines whether the query is complex. If yes, go to step 2, otherwise, set the input query as a single sub-query and go to stage 2 directly.
  2. (2)
    Query decomposer generates initial sub-queries from the input query.
After we have all the sub-queries from stage 1, the system will generate and execute the operators from these sub-queries, as shown in Figure 2 stage 2:
  1. (1)
    For the current sub-query, Meta Engine generates a ranked list of possible operators to be processed with the data of related modality. The details are in Section 4.2.
  2. (2)
    The operators in the list will be routed to proper adapters in the order of rankings and executed one by one, until the first valid and meaningful result $R$ is returned from the execution. Then $R$ is the result for the current sub-query. Additionally, $R$ will be used to refine next sub-query, as introduced in Section 4.2.
  3. (3)
    Repeat Step 1 and 2 until all sub-queries are processed and their results $R_{1},R_{2},R_{3},$... are collected.
After we collected all sub-queries’ answers, the system will put everything into a final answer aggregator, as shown in Figure 2 stage 3, which will comprehend the results into the final result.
Such a workflow ensures (1) the complex query can be analyzed more accurately as single-hop sub-queries are more tractable by the LLM-based semantic systems; (2) the sub-queries are progressively refined with previous sub-query results, reducing sub-query ambiguity; (3) the final result is concise enough after the final aggregation. In summary, it effectively guarantees the result quality.
3.3. Operators
3.3.1. TextAnalytics
operator processes sub-queries whose supporting evidence resides in textual documents. Given a decomposed question and its associated passages, it extract relevant facts, entities, or relations and produces a modality-specific answer conditioned solely on the provided text.
3.3.2. TableAnalytics
operator is designed for processing sub-queries grounded in structured tabular data. It executes operations such as filtering, row/column selection, aggregation, comparison, and numerical computation according to the query intent through assigned adapters.
3.3.3. ImageAnalytics
operator answers sub-queries that require visual evidence from images, charts, or diagrams. It understands semantic information, including objects, labels, tr ends, and numeric values, and aligns these signals with the sub-query through multimodal reasoning.

### v2

# arXiv:2602.01701v2[cs.DB] 02 Mar 2026

@@ Abstract.
Many LLM-based semantic query systems have been proposed to support semantic querying over structured and unstructured data. However, this rapid growth has produced a fragmented ecosystem. Applications face significant integration challenges due to (1) disparate APIs of different semantic query systems and (2) a fundamental trade-off between specialization and generality. Many semantic query systems are highly specialized (e.g., DocETL for long documents, StructGPT for table QA), offering state-of-the-art performance within a specific task or a single modality but struggling with multi-task or multi-modal data. Conversely, some ”all-in-one” systems (e.g., LOTUS) handle multiple modalities but often exhibit suboptimal performance compared to their specialized counterparts in specific modalities. In short, the existing systems still have significant limitations when applied to comprehensive tasks over complex multi-modal data.
This paper introduces Meta Engine, a novel framework for building ”query system on query systems”, designed to resolve those aforementioned challenges. Meta Engine is a unified framework to integrate heterogeneous, specialized LLM-based query systems into one query processing pipeline, such that the specialities of different systems are combined to handle complex scenarios. It also provides unified interfaces for users to query through the pipeline, keeping the disparate APIs of underlying systems transparent to users. Meta Engine integrates several key components for APIs, query preprocessing, query planning and execution, as well as postprocessing, which cover the whole semantic querying lifecycle. The components are extensible, flexible, and composable. Users can extend or replace any of them with their own implementation, and flexibly choose components to customize their pipelines. This design makes Meta Engine easy to use and able to combine the distinct strengths of different semantic query systems, achieving high-performance multi-modal semantic processing from a unified entry point. In the evaluation, Meta Engine consistently outperforms all baselines, yielding 3–6x higher F1 in most cases and up to  24x on specific datasets.
@@ 1. Introduction
Semantic query has become more and more demanded in data management systems, which is an important way to access and analyze unstructured data from various modalities. Comparing to traditional SQL queries that use standard and structured syntax to query the data, semantic queries usually rely on natural language (NL) to semantically process unstructured data like image, text, etc. Therefore, semantic query processing requires the capabilities of understanding NL request and capturing the semantics in data, which is hard in the past years and the solutions are usually limited (Lu et al., 2002; Spanos et al., 2012; Cappuzzo et al., 2020; Toenshoff et al., 2023; Bandyopadhyay et al., 2020). The advance of Large Language Models (LLMs) leads to a paradigm shift in this, resulting in the development of advanced semantic query systems (Cappuzzo et al., 2020; Toenshoff et al., 2023; Bandyopadhyay et al., 2020; Patel et al., 2024; Shankar et al., 2024; Jiang et al., 2023; Khattab et al., 2023; Zhang et al., 2025). They enable applications to semantically reason and analyze over complex (like multi-modal) datasets using natural language.
Despite this rapid progress, the current ecosystem of LLM-based query systems is highly fragmented, presenting critical challenges that hinder the comprehensive use of them on complex tasks. First, diverse systems result in disparate and incompatible APIs. Consequently, developers seeking to build applications that leverage the strengths of multiple systems are forced to write complex integration logic, increasing both development and maintenance overhead. Second challenge is the crucial trade-off between specialization and generality. Many state-of-the-art semantic query systems are highly specialized, achieving exceptional performance by optimizing for specific tasks in single modality. For instance, DocETL (Shankar et al., 2024) excels at processing large-scale long text documents, while StructGPT (Jiang et al., 2023) is designed specifically for semantic querying of tabular data. Their high specialization in specific domain/modality prevents them from integrating and processing information across different modalities. Conversely, ”all-in-one” systems like LOTUS (Patel et al., 2024) attempt to provide a unified solution for multi-modal data (e.g., table, text, and image), but the cost of such a generalization is often suboptimal performance on specific single modality, like the image QA quality of LOTUS is usually lower than a pure VLM in our evaluation.
@@
In this paper, we propose Meta Engine, a novel framework to build the ”query system over query systems”. Meta Engine functions as an comprehensive framework with flexible, extensible and composable components, designed for integrating heterogeneous LLM-based semantic query systems into a cohesive pipeline. By adaptively scheduling the different systems, it can intelligently combine their specialities to achieve a significant performance gain on complex analytics. Additionally, to address the disparate API issue, Meta Engine provides two sets of unified interfaces: a natural language interface and a set of functions including data operators and building block operators. With the interfaces, users can customize the pipeline flexibly, and query the pipeline using either NL questions or programmatically calling the specific operators. These interfaces keep the disparate APIs of underlying systems transparent to users, as well as ensure an easy pipeline construction. Furthermore, Meta Engine supports pipeline construction for not only typical question answering, but also broader data analytic tasks on multi-modal data. For example, in one of our evaluation dataset  (Rahman et al., 2025), each natural language query specifies an analytical task over structured tables, text and images. An example query and supporting documents are shown in Figure 1.
@@
Meta Engine includes several major components: (1) APIs: The APIs include an NL query interface and several functions for the data operators and the building blocks of the pipeline. Table 1 lists the major APIs. Specifically, the building blocks are those used to build the pipeline, while the data operators are those doing the analytics over data in runtime. (2) Query preprocessing: We design the Query Decomposer that smartly breaks down complex NL queries into simpler sub-queries. This component accepts the user’s NL queries and intelligently decompose each query to several sub-queries. A dependency graph is used to manage the dependencies between subqueries for each query. (3) Query planning: After decomposition, Meta Engine maps each subquery to an executable data operator. Since each sub-query is simple enough, one analytic data operator can process it. Instead of selecting a single operator, Meta Engine will select multiple candidate data operators ranked by the selection confidence, where higher-ranked operators are more likely to correctly process the sub-query. More details are in Section 3.1.2 and 4.2. (4) Query execution: Query execution includes execution of data operators and subqueries. Essentially, executing a subquery is to execute its corresponding data operators, with additional processing to the subquery afterwards. In data operator level, an operator will be dispatched to the most appropriate specialized system (adapter) by a Query Router to execute. In subquery level, for each subquery, its ranked candidate operators (from query planning) will be executed one by one as above, following the confidence order (from high to low), until any one successfully processes the subquery and returns valid result. Furthermore, in query level, we propose two optimizations: iterative sub-query refinement and dependency-based subquery parallelism for the execution. More details are in Section 4.2. (5) Adapters: As an important part of query execution, adapters serve as wrappers for the underlying LLM query systems to abstract away the heterogeneity of those systems. Each adapter manages the unique interfaces of a system and exposes standardized interfaces to Meta Engine users. (6) Query postprocessing: After all subqueries are executed, Meta Engine provides an optional Result Aggregator to comprehend the results of sub-queries into the coherent final result, with necessary refinement like deduplicating. If users tend not to do the postprocessing, the final result will be the result of the last subquery.
Additionally, Meta Engine is extensible that allows users to extend most components, e.g., users can replace the decomposer with their own implementation. Meta Engine is also composable and flexible, as most components are optional in a pipeline and users are free to build the pipeline as they like. For example, users can pass the original query to select_opselect\_op without decomposition, or directly call specific data operator to process the query/subquery while skipping select_opselect\_op. With the composable, extensible components and unified APIs in Meta Engine, users can easily customize a pipeline that integrates different specialized systems for complex multi-modal analytics. The architecture and optimization in Meta Engine resolves the specialist-generalist trade-off, enabling applications to benefit from high-performance, multi-modal processing with a unified solution. To our best knowledge, Meta Engine is among the first works to build such a unified and flexible framework for integrating heterogeneous LLM systems in one pipeline.
@@
    We propose Meta Engine, a novel unified framework to efficiently build multi-modal query processing pipeline that integrates different specialized LLM query systems.
@@
    We develop composable, extensible and flexible architecture and components, with novel optimization techniques.
@@
    We design unified APIs to unify the heterogeneous interfaces of different query systems, significantly simplifying pipeline development.
@@
    We conduct extensive evaluation that shows the effectiveness of our proposed system, as well as ablation studies to show the impacts of each major component.
The remainder of this paper is structured as follows: Section 2 introduces prior works related to Meta Engine. Section 3 and 4 introduce the architecture, workflow and technical details of our method. Section 5 reports and analyzes experiment results.
Table 1. Summary of Key APIs: (1) End-to-end NL interface, (2) building block APIs for pipeline customization, and (3) data operator APIs for multi-modal analytics.
API Return Description Algorithm
End-to-End Interface
query($q$) Query result End-to-end pipeline execution for NL query $q$.
Building Block APIs
decompose($q$, checkcheck, max_outmax\_out) Sub-queries Decomposes $q$ into sub-queries; conditional on complexity if $check{=}\text{True}$. Dynamic decomposition
select_op($q$, opsops) Ranked op. list Ranks candidate operators for $q$ by confidence. LM-based classification
exec_op(opsops, rr:{S||L}) Op. results Routes operators to adapters and executes; $r$ selects routing strategy. History / trained routing
exec_subq(qsqs, parpar) Subquery results Executes sub-queries via select_op and exec_op; parpar enables parallelism. Iterative refinement + parallel exec.
route(opsops, adptsadpts, rr:{S||L}) Adapter assign. Maps each operator to its best adapter; $r$ selects routing strategy. Statistics / learning-based pred.
agg_results(resres, $q$) Final answer Aggregates sub-query results into the final answer for $q$. LLM summarization
Data Operator APIs
TextAnalytics($q$) Text analytics result Executes analytics specified by $q$ on text documents. Adapter-based (text)
ImageAnalytics($q$) Image analytics result Executes analytics specified by $q$ on image documents. Adapter-based (image)
TableAnalytics($q$) Table analytics result Executes analytics specified by $q$ on tables. Adapter-based (table)
@@ 3. Architecture
Figure 2. Architecture of Meta Engine
Figure 2 presents the architecture and query/data flow of Meta Engine.
@@ 3.1. Architecture
Meta Engine has a modular architecture that coordinates the processing of complex, multi-modal queries through specialized components. As shown in Figure 2, the architecture includes three levels: Top-level APIs (red box), middle-level building blocks (blue boxes), and bottom-level adapters (green box). Users use APIs to interact with building blocks, building blocks work with each other and call adapters to process the query, while adapters access and operate the data.
3.1.1. APIs
As listed in Table 1, Meta Engine provides three types of APIs: (1) End-to-end query interface query($q:Str$), which accepts the users’ natural language query and returns the final result; (2) Building block interfaces, which allows users to access the key components in Meta Engine for building and customizing their pipelines, ranging from query decomposer to result refiner; (3) Data operator interfaces, which call the adapters to execute specific analytics over data in runtime. The APIs are composable and flexible that users can freely choose and combine some of them to construct a semantic query pipeline on top of multiple specialized query systems. And all APIs are extensible, i.e., Meta Engine allows users to easily replace or modify the implementation of the APIs. Such APIs significantly save users’ development cost, enabling efficient construction of semantic analytics pipeline on multi-modal data.
3.1.2. Building Blocks
The building blocks are categorized by four stages of the query processing: query preprocessing, query planning, query execution and query postprocessing, indicated by the blue boxes in Figure 2. The dashed boarder boxes represent the stages while the solid boarder boxes inside each stage present the components working for the stage.
The first stage is query preprocessing (first blue box in Figure 2) which checks and decomposes the input query into several subqueries. The component in this stage is the Query Decomposer. We design a dynamic decomposition algorithm that decomposes each query based on its complexity. Specifically, if the query is multi-hop, it will be decomposed to several single-hop subqueries, while if it is already single-hop, decomposition will not be conducted. The API for calling query decomposer is $\texttt{decompose}($q:Str,:Str,check:Bool,:Bool,max_out:Int):Int), which accepts an NL query $q$ and decomposes it into simpler sub-queries. Parameter checkcheck specifies whether to enable the dynamic decomposition: if $check=True$, the query may or may not be decomposed, depending on its complexity; if $check=False$, it is forced to always decompose the query, even if it is already single-hop. max_outmax\_out specifies the maximum number of subqueries the users want. If the decomposer generates more subqueries, it will be required to re-do the decomposition with emphasis on the number of subqueries. More details are described in Section 4.1.
The second stage is query planning (second blue box in Figure 2), which maps the NL subqueries to data operators. In our current implementation, each subquery will be mapped to one data operator among the three, i.e., TextAnalytics, ImageAnalytics, and TableAnalytics. Because each subquery is simple and single-hop, one data operator is enough to process it. The core component in this stage is Operator Generator, responsible for the Operator Selection, i.e., identifying and selecting the most appropriate analytical operators for each subquery. This mirrors the logical plan generation in DBMS, i.e., from query to logical operator(s). A key challenge in multi-modal data analytics is that it’s unknown which modalities contain the necessary information to the query/subquery. Even though each subquery is simple, it is still hard to predict which modality of data is related to this subquery. Accessing all modalities could ensure the needed information is covered, but this solution is computation-intensive. Predicting just the most likely modality is more efficient, but wrong prediction will lead to meaningless result. To achieve maximum information coverage while maintaining good efficiency, we design a ranking-based operator selection and execution mechanism. Specifically, the operator generator will return a ranked list of multiple candidate data operators, rather than only selecting one operator. Then the candidate operators will be executed progressively with early stopping in execution stage. More details are introduced in Section 4.2. The API to call the operator generator is select_op($q$: Str, opsops: List). It accepts query $q$ and a list of available data operators (all the three data operators in Meta Engine by default), returns a ranked list of candidate operators, ordered by the determination confidence. This is completed by a language model based classification.
The third stage is query execution (the third blue box in Figure 2). As an important stage in Meta Engine, query execution comprises of several components: the Query Router and various Adapters. An adapter is a wrapper of an existing semantic query system, such as LOTUS, DocETL, StructGPT, DSPy and so on. On one hand, these adapters encapsulate the disparate interfaces of various systems and provide unified data operator APIs for users and other components in Meta Engine. This significantly saves development cost. On another hand, adapters directly interact with the multi-modal data sources (tables, text, and images) to perform the operator/subquery, ensuring the power of different systems to be jointly utilized.
The query router sends each sub-query’s data operator to the proper adapter to execute. This step mirrors the physical plan generation in DBMS, i.e., from logical operator to physical implementation, and here the physical implementation is the concrete implementation of the data operator in each underlying system. Note that each data operator is associated with the corresponding subquery (like TextAnalytics($q$: Str)), so the router makes the decision based on both the operator type and the subquery. Furthermore, we design two types of routers, statistics-based and learning-based, with different advantages and satisfying different demands. More details are in Section 4.4.
Additionally, we propose novel optimizations for subquery execution, including iterative subquery refinement and dependency-based parallelism, which progressively refine the next subquery with previous results, and parallelize the subquery execution based on their dependencies. More details are in Section 4.2.
There are several APIs associated with this execution stage: exec_subq($qs$: List, parpar: Bool), exec_op(opsops: List, $r$: {S||L}), and route(opsops: List, adptsadpts: List, $r$: {S||L}). These APIs are implemented from top to bottom. The top level is exec_subq that utilizes select_op and exec_op to do logical planning, physical planning and execution for each given subquery. Parameter parpar determines whether to enable the dependency-based parallelism for subqueries. The middle level is exec_op, which uses route to route each data operator in opsops to an adapter, call the adapter to execute it, and return the results. At the bottom, route predicts a proper adapter per data operator. Parameter $r$ specifies which routing algorithm to use, Statistics-based prediction ($S$) or learning-based classification ($L$).
Finally, in the postprocessing stage (the fourth blue box in Figure 2), the partial results from these isolated sub-query executions are synthesized by the Result Aggregator. This component cleans the results from all sub-queries and comprehends them to produce the final result to the original user-input query. The API to call this aggregator is agg_results(resres: List, $q$: Str). It aggregates and refines the subquery results in the list resres into the final result to query $q$. An LLM will be used to conduct the aggregation. If a pipeline does not use the aggregator, the final result will be the result of the last subquery. More details are in Section 4.3.
With such an architecture, Meta Engine demonstrates a significant improvement in result quality compared to the state-of-the-art LLM-based query systems. By the joint use of specialized LLM systems and novel optimizations, Meta Engine consistently achieves higher accuracy with competitive efficiency in our evaluation. Another core advantage of the architecture is its modular design. The components are flexible to be composed into customized pipelines. Each component has an API for users to directly access. The underlying LLM systems are wrapped by adapters and transparent to users. This design decouples the reasoning logic from the backend execution, allowing for seamless addition or removal of underlying systems (e.g., LOTUS, DSPy, DocETL) without overhauling the core engine, and enabling flexible pipeline construction. With new powerful LLM systems emerge, they can be easily integrated as new adapters. We demonstrate several different pipelines built upon Meta Engine in Section 3.2.
3.1.3. Data Operators and Adapters
Currently Meta Engine has three data operators, shown in Table 1: (1) TextAnalytics operator (API TextAnalytics($q$: Str)) conducts analytics whose supporting evidence resides in textual data. Given a query/subquery, this operator extracts relevant facts, entities, or relations and produces the analytic results conditioned solely on the text data. (2) TableAnalytics operator (API TableAnalytics($q$: Str)) is designed for processing analytics grounded in structured tabular data. Its processing includes filtering, row/column selection, aggregation, comparison, and numerical computation according to the query intent. (3) ImageAnalytics operator (API ImageAnalytics($q$: Str)) processes visual analytics that requires visual evidence from images, charts, or diagrams. It understands semantic information, including objects, labels, trends, and numeric values in the figures, and aligns these signals with the sub-query to generate results.
Each data operator has an execute interface in one or more adapters, which acts as the physical implementation of this operator. An adapter may not implement the interfaces for all data operators, depending on its speciality and limitation. For example, in our current implementation, StructGPT adapter only has execute interface for TextAnalytics, as StructGPT is designed to process structured data; while LOTUS, DSPy and Single-Model adapters support all three data operators. As a result, a data operator could be supported by multiple adapters, like LOTUS, DSPy and Single-Model adapters all support ImageAnalytics. In runtime the router will choose the potentially best adapter to execute the data operator, i.e., predicting and selecting the adapter which is likely to have the best performance to execute the given data operator.
Figure 3. Example pipelines built with Meta Engine framework
3.2. Pipeline Construction
Users are free to use different components of Meta Engine to customize their pipelines. Figure 3 demonstrates some example pipelines that can be built with Meta Engine framework. To make the figures clear, we ignore the boxes of APIs and adapters. We use “Subq”, “Op”, “Q/Subq” to denote “Subqueries”, “Operators” and “Query/Subqueries”.
Pipeline 1 is a fully automated pipeline with all the major components introduced above, from query decomposer to result aggregator. The input is the original user query, going through the whole pipeline until the final result is made. Such a pipeline provides the easiest way to use. Pipeline 2 is a partially automated pipeline without pre and post processing, which is for the scenarios that users have subqueries ready or prefer to keeping the query as a whole. This pipeline saves the time spent on decomposition and aggregation. Pipeline 3 is a mini pipeline with the minimum number of components, which gives users the most control of the query processing. Users can prepare the query/subqueries in their own way and customize the operator assignment for the query/subqueries. This pipeline only takes responsibility of executing the operators. Such a pipeline is useful in some cases, e.g., the case that users know the needed data modality. Like in a text QA task, the only needed data operator is TextAnalytics, thus the operator selection can be skipped to save time and tokens. In extreme cases, the router can be removed if users prefer a specific underlying system.
In summary, Meta Engine framework is highly composable and flexible. Users can build an end-to-end pipeline including all stages or customize a pipeline that starts from a specific stage. The query preprocessing and postprocessing stages are always optional in a pipeline. Without decomposer, users can input the original query or prepared subqueries to the pipeline, while without result aggregator, the last subquery’s result will be considered as the final result. Query planning stage (operator selection) can also be ignored if users prefer manually setting the data operators. The only required stage in a pipeline is the execution, including query routing and execution by adapter. Such a framework provides the maximum flexibility to pipeline construction for different demands. In addition, Meta Engine also allows the pipelines to be controllable. Since each component is extensible, users can easily control the number and behavior of the components for pipeline diagnosis. For example, users can control the router to send all subqueries to a specific adapter by purpose, to evaluate the underlying system’s performance in the pipeline.